# Quantum metrology through spectral measurements in quantum optics

### v1

@@ I Introduction
Figure 1: (a) Sketch of the metrological setup. A coherently driven emitter, characterized by atomic parameters to be estimated, $\theta\in\{\Delta,\Omega,\gamma\}$, whose output field is divided by a beam splitter and subsequently filtered by two frequency-selective filters centered at $\Delta_{i}$ with spectral resolution $\Gamma_{i}$. (b) Metrological protocol. A sensor collects the radiation emitted by the source, giving access to the filtered density matrix of the light in the frequency domain. This process allows the extraction of the $\theta$-dependent photon-number distribution, $p(n|\theta)=\langle n|\hat{\rho}_{\theta}|n\rangle$ (top panel). An estimation strategy can then be employed to infer $\theta$ from $p(n|\theta)$; the precision of any unbiased estimator is bounded by the inverse of the Fisher information (bottom panel). The solid black curve represents the quantum Fisher information ($I_{\Delta}$) and the solid blue curve the classical Fisher information ($F_{\Delta}$). The dashed blue curve the signal-to-noise ratio ($\text{SNR}_{\Delta}$), which gives the precision of an estimation based only on the mean population of the sensor $\langle\hat{\xi}^{\dagger}\hat{\xi}\rangle$ (black-dashed curve in the top panel) instead of the full photon-counting distribution. All quantities correspond to the inference of the qubit–laser detuning parameter $\Delta$. (c) Joint $\theta$-dependent probability distribution, $p(n_{1},n_{2}|\theta)$ for simultaneous detection with two sensors. The top and right panels display the marginal distributions for each detector, $p(n_{i}|\theta)$, highlighting their sensitivity to a small perturbation in the parameter, $\theta+\delta\theta$ (a non-infinitesimal perturbation $\delta\theta=-0.65\gamma$ was chosen to enhance the discernibility of the curves). Parameters: (b) $\Omega=10\gamma$, $\Delta=10^{-5}\gamma$, $\Gamma=10^{-1}\gamma$, $\varepsilon=0.9$. (c) $\Omega=2\gamma$, $\Delta=0$, $\Gamma=10^{-2}\gamma$, $\varepsilon=0.75$, $\Delta_{1}=0$, $\Delta_{2}=5\cdot 10^{-3}\gamma$;
@@
Continuously monitored open quantum systems are emerging as promising platforms for quantum metrology [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], with key applications in areas such as atomic magnetometry [19, 20], force sensing [21], gravitational wave detection [22], and criticality-enhanced metrology [23, 15, 24, 25, 26, 27, 28, 29, 30, 31, 16, 32, 33]. These developments have been accompanied by significant theoretical progress, particularly in establishing a formal framework for quantum metrology in such systems. Notably, this includes the derivation of ultimate precision bounds based on the Cramér–Rao bound for information encoded in the radiated field [34, 35, 24].
@@
However, a major open challenge remains: devising measurement strategies that efficiently extract the full information content from the high-dimensional quantum state of the radiated field [4]. Standard time-local measurements, such as photon-counting or homodyne detection, fall short of saturating the quantum Fisher information (QFI) [34, 36], as they fail to capture information embedded in higher-order and non-time-local correlations. To overcome this limitation, recent strategies go beyond time-local measurements by incorporating ancillary systems—such as absorbers [37] or decoders [4]—that interact with and extract structured information from the radiated field. These approaches are closely tied to efforts aimed at reconstructing the full quantum state of the output field, often by coupling to ancillary modes designed to capture specific temporal profiles [38, 39, 40]. This filtering of the output field naturally connects with frequency filtering, which can be understood as a physically motivated temporal-mode basis [41] which, notably, it can also be formally described in terms of ancillary modes [42, 43, 44]. Frequency filtering is routinely implemented in experimental setups via interferometers and spectrometers [45, 46, 47, 48], providing a direct and accessible way to probe structured features of the radiated field.
@@
Despite this conceptual alignment, frequency-resolved measurements have yet to be fully leveraged in the context of quantum metrology. Although multi-photon frequency correlations have been extensively studied—particularly in systems such as coherently driven two-level emitters [47, 45, 49, 50, 51, 52, 53, 40], cavity QED [49], optomechanical [54] and molecular platforms [55, 56]—their potential as a resource for parameter estimation and precision sensing remains largely untapped. This represents a significant opportunity, as the light emitted by continuously driven quantum systems displays rich structures in their frequency-frequency correlation functions which are expected to encode valuable information about the underlying system parameters.
@@
We apply our formalism to a minimal yet highly instructive system: a single two-level emitter coherently driven near resonance. Despite its simplicity, this setup exhibits considerable complexity and reveals a rich landscape of multi-photon processes [42, 47, 45, 49, 50, 51, 52, 53, 40]. It serves as an ideal testbed for exploring the emergence of non-classical correlations in the emitted light and for addressing some of the key questions outlined above, exploring the role of such correlations in enhancing quantum parameter estimation of unknown atomic parameters.
We assess the metrological potential of frequency filtering by evaluating the classical Fisher information associated with coherently displaced photon-counting measurements of frequency filtered modes. Specifically, we quantify the sensitivity of estimating an unknown atomic parameter based on photons detected at a given frequency $\omega$, with spectral resolution characterized by a linewidth $\Gamma$, and analyze how this estimation is enhanced when frequency correlations between photons detected at different frequencies are taken into account.
The manuscript is organized as follows. In Section II, we introduce the model and outline the metrological strategy. In Section III, we focus on the single-sensor case, where we define the frequency-resolved Fisher information and analyze its key features, including its spectral dependence and sensitivity to the filtering process. We also present a mean-field engineering approach to optimize the Fisher information. Finally, in Section IV, we extend the framework to two sensors, demonstrating how photon-photon correlations can be exploited to improve parameter estimation.
@@ II General framework
In this section, we introduce the physical model and the metrological strategy for estimating unknown atomic parameters. While the framework is fully general and applicable to a broad class of systems, we focus here on a minimal implementation sketched in Fig.1(a): frequency-filtering of the radiation emitted by coherently driven quantum emitter. This setting allows us to clearly illustrate the core features and advantages of the proposed method.
@@ II.1 Model
We describe the quantum source as a coherently driven two-level system (TLS), spanning a basis $\{|g\rangle,|e\rangle\}$, which denotes the ground and excited states, respectively, such that the lowering operator gets defined as $\hat{\sigma}\equiv|g\rangle\langle e|$. The quantum emitter is driven by a coherent field characterized by a Rabi frequency $\Omega$ and a laser frequency $\omega_{L}$. In the rotating frame of the laser, and under the rotating wave approximation, the Hamiltonian of the source reads
@@
where $\Delta\equiv\omega_{\sigma}-\omega_{L}$ is the qubit-laser detuning, with $\omega_{\sigma}$ being the natural frequency of the quantum emitter. The source undergoes incoherent processes, essentially de-excitation through spontaneous emission at a rate $\gamma$, resulting from its interaction with the vacuum electromagnetic field or a structured environment such as a coplanar waveguide [57, 58, 59]. The dissipative dynamics of the TLS, described by its reduced density matrix $\hat{\rho}$, is governed by the following master equation:
$\frac{d\hat{\rho}}{dt}=-i[\hat{H},\hat{\rho}]+\frac{\gamma}{2}\mathcal{D}[\hat{\sigma}]\hat{\rho}$, (2)
@@
In order to describe the quantum state of specific frequency-filtered modes of the emitted field, we will recourse to the cascaded-sensor method [42, 43, 44]. The corresponding master equation is given, in the case of a single sensor, by [see details in Appendix A]
@@
where the sensor is described by a single bosonic mode with annihilation operator $\hat{\xi}$, sensor-laser detuning $\Delta_{\xi}\equiv\omega_{\xi}-\omega_{L}$ ($\omega_{\xi}$ being the sensor frequency), and linewidth $\Gamma$. Additionally, we have introduced an imperfect coupling factor $\varepsilon\leq 1$, accounting, e.g., for possible losses in the transmission from the source to the target [60], or to retain the scattered light from a coherently driven system [43, 44]. Therefore, $\varepsilon=1$ corresponds to perfect cascaded coupling, and $\varepsilon=0$ to the limit of no coupling.
Although the single-sensor case is relatively straightforward, the extension to two or more cascaded sensors requires additional attention, since there are multiple descriptions which correspond to different physical implementations of the multi-mode measurement. In this work, we will explore the specific case in which, prior to the filtering of each mode, the output signal is physically split [61, 53] (e.g. by a balanced beam splitter in a Hanbury-Brown Twiss setup), as illustrated in Fig. 1(a). The master equation now takes the form [see details in Appendix A]
$\frac{d\hat{\rho}}{dt}=-i[\hat{H}+\Delta_{1}\hat{\xi}_{1}^{\dagger}\hat{\xi}_{1}+\Delta_{2}\hat{\xi}_{2}^{\dagger}\hat{\xi}_{2},\hat{\rho}]+\frac{\gamma}{2}\mathcal{D}[\hat{\sigma}]\hat{\rho}\\ +\sum_{i=1}^{2}\frac{\Gamma_{i}}{2}\mathcal{D}[\hat{\xi}_{i}]\hat{\rho}-\sqrt{\varepsilon\gamma\Gamma_{i}/2}\left\{[\hat{\xi}_{i}^{\dagger},\hat{\sigma}\hat{\rho}]+[\hat{\rho}\hat{\sigma}^{\dagger},\hat{\xi}_{i}]\right\}.$ (4)
In both cascaded master equations, Eq (3) and Eq. (4), $\hat{H}$ denotes the Hamiltonian of the quantum source as given in Eq. (1). The operator $\hat{\xi}_{i}$ represents the $i$th bosonic mode associated with each sensor, characterized by a sensor-laser detuning $\Delta_{i}\equiv\omega_{i}-\omega_{L}$, and linewidth $\Gamma_{i}$. To relax the notation in the two-sensor case, we have replaced the variable $\Delta_{\xi}\rightarrow\Delta_{i}$ ($i=1,2$). Throughout this work we consider identical sensor linewidths, $\Gamma_{1}=\Gamma_{2}=\Gamma$. The factor $1/\sqrt{2}$ in the cascaded-coupling term of Eq. (4) accounts for vacuum contributions arising from signal splitting at the balanced beam splitter. In contrast, in a previous work [40], we considered the situation in a waveguide setup in which the photonic modes are digitally filtered by post-processing the signal, resulting in a different master equation from Eq. (4) that does not include the factor $1/\sqrt{2}$ in the cascaded-coupling term.
@@ II.2 Metrological strategy
where $\hat{D}(\vec{\alpha})=\prod_{i=1}^{N}\hat{D}(\alpha_{i})$ is the multi-mode displacement operator, with each $\alpha_{i}\in\mathbb{C}$ corresponding to the coherent displacement applied to the $i$th sensor, and $|\vec{n}\rangle\equiv|n_{1},\ldots,n_{N}\rangle$, with $|n_{i}\rangle$ denoting the photon number eigenstates of the $i$th sensor. We implement this strategy by performing $N$ discrete frequency-resolved displaced photon-counting measurements, with each frequency channel modeled as a single-mode bosonic cavity. Formally, the measurement process is described by the tensor product of sets of operators describing the POVMs, $\hat{\Lambda}=\bigotimes_{i=1}^{N}\hat{\Lambda}_{n,\alpha_{i}}^{(i)}$, where
$\hat{\Lambda}_{n,\alpha_{i}}^{(i)}=\{\hat{D}^{\dagger}(\alpha_{i})|n_{i}\rangle\langle n_{i}|\hat{D}(\alpha_{i})\}.$ (7)
Physically, displaced photon-counting can be realized by mixing the output field of the quantum emitter with a local oscillator $\alpha(t)$ via a highly transmissive beam splitter, yielding an effective detection signal $\sqrt{\gamma}\hat{\sigma}(t)+\alpha(t)$. This mean-field engineering technique [71, 72] enables us to tailor the photonic state to either isolate the quantum fluctuations or optimize the entire state for the estimation of specific atomic parameters. Standard photon-counting corresponds to the special case $\vec{\alpha}=\vec{0}$, denoted as $p(\vec{n}|\theta)\equiv p^{\vec{\alpha}=\vec{0}}(\vec{n}|\theta)$ and $F_{\theta}\equiv F^{\vec{\alpha}=\vec{0}}_{\theta}$.
Although this formulation applies generally to any number of frequency-resolved sensors, we restrict our analysis to the one‐ and two‐sensor cases. Accordingly, we evaluate $p^{\alpha}(n|\theta)$ for $N=1$ and $p^{(\alpha_{1},\alpha_{2})}(n_{1},n_{2}|\theta)$ for $N=2$, as illustrated in Fig. 1(b) and (c), respectively. In this work, $\theta$ represents a Hamiltonian or a Lindblad parameter from the source master equation in Eq. (2), such as the qubit-laser detuning, the driving strength of the laser interacting with the quantum emitter, or the emitter decay rate, $\theta\in\{\Delta,\Omega,\gamma\}$.
@@
$\Delta^{2}\hat{\theta}\geq\frac{1}{MI_{\theta}}\geq\frac{1}{MF_{\theta}^{\vec{\alpha}}}\geq\frac{1}{M\text{SNR}_{\theta}[\hat{O}]}$. (9)
where $M$ is the number of repetitions, and $I_{\theta}$ denotes the quantum Fisher information (QFI), defined as the optimization of the CFI over all POVMs, or equivalently, the optimization of the SNR over all observables [17]. Throughout this work, we use these three figures of merit—QFI, CFI, and SNR—to characterize the metrological performance of our proposed strategy. We finally give here a relevant caveat regarding the QFI. Here, the QFI is computed over the quantum state of the sensors, which captures the filtered version of the emitted light. While it bounds the precision achievable with optimal measurements on the sensor modes, it does not necessarily correspond to the absolute lower bound for $\Delta^{2}\hat{\theta}$, since it excludes information potentially available in the total radiation of the emitter, quantified by general expressions of the QFI as developed in Refs. [34, 24].
Figure 1(b) illustrates the application of this framework for the single-sensor case. The radiation emitted by the source is filtered by the sensor, yielding a frequency-resolved density matrix from which the conditional photon-count distribution $p(n|\theta)$ is extracted [top panel]. In this example, the distribution, obtained via standard photon-counting measurements ($\alpha=0$), forms the basis for estimating the parameter $\theta$ with a precision set by the CFI, shown by the blue solid curve. The example highlights the advantage of using the full photon-count statistics over relying only on the mean photon number. This is evident from the comparison between the CFI and the SNR of the sensor photon number (blue dashed curve). Both quantities lie below the QFI (black solid curve), indicating that neither corresponds to the optimal measurement for estimating the detuning ($\theta=\Delta$). As an illustration of the extension to the two-sensor case, Fig. 1(c) displays the joint distribution $p(n_{1},n_{2}|\theta)$ for $\alpha=0$ for the detection at two different frequencies, revealing off-diagonal structure that signals cross-correlations between detection channels, a feature that can be harnessed to enhance parameter estimation.

### v2

@@ I Introduction
Figure 1: (a) Sketch of the metrological setup. A coherently driven emitter—characterized by the parameter to be estimated, $\theta\in\{\Delta,\Omega,\gamma\}$—emits radiation that is split by a beam splitter and routed into two frequency-selective sensors, each centered at $\Delta_{i}$ and featuring a spectral resolution $\Gamma_{i}$. (b) Metrological protocol. A sensor collects the radiation emitted by the source, giving access to the filtered density matrix of the output field in the frequency domain. This process allows the extraction of the $\theta$-dependent photon-number distribution, $p(n|\theta)=\langle n|\hat{\rho}_{\theta}|n\rangle$ (top panel). An estimation strategy can then be employed to infer $\theta$ from $p(n|\theta)$. The precision of any unbiased estimator is bounded by the inverse of the Fisher information (bottom panel): $I_{\Delta}$ (solid black), $F_{\Delta}$ (solid blue), and $\text{SNR}_{\Delta}$ (dashed blue), which gives the precision of an estimation based only on the mean population of the sensor $\langle\hat{\xi}^{\dagger}\hat{\xi}\rangle$ (black-dashed curve in the top panel) instead of the full photon-counting distribution. All quantities correspond to the estimation of the qubit–laser detuning parameter $\Delta$. (c) Joint $\theta$-dependent probability distribution, $p(n_{1},n_{2}|\theta)$ for simultaneous detection with two sensors. The top and right panels display the marginal distributions for each detector, $p(n_{i}|\theta)$, highlighting their sensitivity to a small perturbation in the parameter, $\theta+\delta\theta$ (a non-infinitesimal perturbation $\delta\theta=-0.65\gamma$ was chosen to enhance the discernibility of the curves). Parameters: (b) $\Omega=10\gamma$, $\Delta=10^{-5}\gamma$, $\Gamma=10^{-1}\gamma$, $\varepsilon=0.9$. (c) $\Omega=2\gamma$, $\Delta=0$, $\Gamma=10^{-2}\gamma$, $\varepsilon=0.75$, $\Delta_{1}=0$, $\Delta_{2}=5\cdot 10^{-3}\gamma$;
@@
Continuously monitored open quantum systems are emerging as promising platforms for quantum metrology [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19], with key applications in areas such as atomic magnetometry [20, 21], force sensing [22], gravitational wave detection [23], and criticality-enhanced metrology [24, 25, 26, 27, 28, 29, 30, 16, 31, 32, 33, 18, 34]. These developments have been accompanied by significant theoretical progress, particularly in establishing a formal framework for quantum metrology in such systems. Notably, this includes the derivation of ultimate precision bounds based on the Cramér–Rao bound for information encoded in the radiated field [35, 36, 24].
@@
However, a major open challenge remains: devising measurement strategies that efficiently extract the full information content from the inherently multi-mode quantum state of the radiated field [14]. Standard time-local measurements, such as photon-counting or homodyne detection, fall short of saturating the quantum Fisher information (QFI) [35, 37], as they fail to capture information embedded in higher-order and non-time-local correlations. To overcome this limitation, recent strategies go beyond time-local measurements by incorporating ancillary systems—such as absorbers [38] or decoders [14]—that interact with and extract structured information from the radiated field. These approaches are closely tied to efforts aimed at reconstructing the full quantum state of the output field, often by coupling to ancillary modes designed to capture specific temporal profiles [39, 40, 41]. This filtering of the output field naturally connects with frequency filtering, which can be understood as a physically motivated temporal-mode basis [42, 43] which, notably, can also be formally described in terms of ancillary modes [44, 45, 46]. The expectation values obtained from the steady state of these ancillary modes correspond, by the ergodic hypothesis, to the time-averaged frequency-filtered signals continuously measured over a single realization [47, 48, 49]. Frequency filtering is routinely implemented in experimental setups via interferometers and spectrometers [50, 51, 52, 53], providing a direct and accessible way to probe structured features of the radiated field.
@@
Despite this conceptual alignment, frequency-resolved measurements have yet to be fully leveraged in the context of quantum metrology. Although multi-photon frequency correlations have been extensively studied—particularly in systems such as coherently driven two-level emitters [50, 51, 54, 55, 56, 57, 58, 41], cavity QED [54], optomechanical [59] and molecular platforms [60, 61]—their potential as a resource for parameter estimation and precision sensing remains largely untapped. This represents a significant opportunity, as the light emitted by continuously driven quantum systems displays rich structures in their frequency-frequency correlation functions, which are expected to encode valuable information about the underlying system parameters.
@@
In most of this work, we apply our formalism to a minimal yet highly instructive system: a single two-level emitter coherently driven near resonance. Despite its simplicity, this setup exhibits considerable complexity and reveals a rich landscape of multi-photon processes [44, 50, 51, 54, 55, 56, 57, 58, 41]. It serves as an ideal testbed for exploring the emergence of non-classical correlations in the emitted light and for addressing some of the key questions outlined above, i.e., exploring the role of such correlations in enhancing quantum parameter estimation of unknown atomic parameters. We assess the metrological potential of frequency filtering by evaluating the classical Fisher information associated with coherently displaced photon-counting measurements of frequency filtered modes. Specifically, we quantify the sensitivity of estimating an unknown atomic parameter based on photons detected at a given frequency $\omega$, with spectral resolution characterized by a linewidth $\Gamma$, and analyze how this estimation is enhanced when frequency correlations between photons detected at different frequencies are taken into account.
The manuscript is organized as follows. In Section II, we introduce the model and outline the metrological strategy. In Section III, we focus on the single-sensor case, where we define the frequency-resolved Fisher information and analyze its key features, including its spectral dependence and sensitivity to the filtering process. We also present a mean-field engineering approach to optimize the Fisher information. In Section IV, we extend the framework to two sensors, demonstrating how photon-photon correlations can be exploited to improve parameter estimation. Finally, in Section V, we illustrate the generality of the framework by analyzing its application to a transmon qubit and an optomechanical system.
@@ II General framework
In this section, we introduce the physical model and the metrological strategy for estimating unknown atomic parameters. While the framework is fully general and applicable to a broad class of systems, we focus here on a minimal implementation sketched in Fig. 1(a): frequency-filtering of the radiation emitted by a coherently driven quantum emitter. This setting allows us to clearly illustrate the core features and advantages of the proposed method.
@@ II.1 Model
We describe the quantum source as a coherently driven two-level system (TLS)—which we will also refer to as a qubit—, spanning a basis $\{|g\rangle,|e\rangle\}$, which denotes the ground and excited states, respectively, such that the lowering operator gets defined as $\hat{\sigma}\equiv|g\rangle\langle e|$. The quantum emitter is driven by a coherent field characterized by a Rabi frequency $\Omega$ and a laser frequency $\omega_{L}$. In the rotating frame of the laser, and under the rotating wave approximation, the Hamiltonian of the source reads
@@
where $\Delta\equiv\omega_{\sigma}-\omega_{L}$ is the qubit-laser detuning, with $\omega_{\sigma}$ being the natural frequency of the quantum emitter. The source undergoes incoherent processes, essentially de-excitation through spontaneous emission at a rate $\gamma$, resulting from its interaction with the vacuum electromagnetic field or a structured environment such as a coplanar waveguide [62, 63, 64]. In the Markovian regime, the dissipative dynamics of the TLS—described by its reduced density matrix $\hat{\rho}_{\mathrm{q}}$—, is governed by the following master equation:
$\frac{d\hat{\rho}_{\mathrm{q}}}{dt}=-i[\hat{H},\hat{\rho}_{\mathrm{q}}]+\frac{\gamma}{2}\mathcal{D}[\hat{\sigma}]\hat{\rho}_{\mathrm{q}},$ (2)
@@
Considering now the output field $\hat{a}_{\text{out}}(t)$ radiated by the source, we can define any temporal mode with bosonic statistics as $\hat{a}_{v}=\int_{-\infty}^{\infty}v^{*}(t)\hat{a}_{\text{out}}(t)\,dt$, provided a normalized filter function, $\int_{-\infty}^{\infty}|v(t)|^{2}\,dt=1$. In order to describe frequency-filtered modes of the emitted field and obtain their full quantum state, we will recourse to the time-independent sensor method [44, 45, 46], which, in the case of frequency-filtering, is completely equivalent to the time-dependent input-output theory of quantum pulses developed in Refs. [39, 40], as we show below. In the latter approach, the system and the emitted radiation are modeled as a cascaded quantum network [48, 65], where the desired temporal mode $\hat{a}_{v}(t)$ is described by an ancillary cavity (with annihilation operator $\hat{\xi}$), driven non-reciprocally by the system with a time-dependent coupling rate
$g_{v}(t)=-\frac{v^{*}(t)}{\sqrt{\int_{0}^{t}dt^{\prime}|v(t)|^{2}}}.$ (3)
It can be shown [39, 40] that this choice of $g_{v}(t)$ results in $\hat{\xi}=\hat{a}_{v}$, meaning that the state of the ancillary sensor matches the state of the selected mode of the radiated field. To describe the particular case of frequency-filtering, we set the temporal mode $v(t)$ as the normalized Lorentzian band-pass filter [42],
$v(t)=\sqrt{\Gamma}e^{(t_{0}-t)(i\omega_{\xi}-\Gamma/2)}e^{i\phi}\Theta(t_{0}-t),$ (4)
where $\omega_{\xi}$ is the frequency of the filtered mode, $\Gamma$ is the linewidth of the Lorentzian filter, $\phi$ is a constant phase, and $\Theta(t)$ is the Heaviside step function. This filter function describes the time-dependent physical spectrum of light [42], with the time-dependence encoded in the free parameter $t_{0}$. Since Eq. (3) assumes $t=0$ as the starting point of the dynamics (with the ancillary modes in vacuum), $v(t)$ is a well-normalized function in the steady-state limit $t_{0}\rightarrow\infty$. In that limit, the time-dependent coupling (in the rotating frame of the drive) simplifies to
$g_{v}(t)=-\sqrt{\Gamma}e^{i(\Delta_{\xi}t+\phi)},$ (5)
where $\Delta_{\xi}\equiv\omega_{\xi}-\omega_{L}$ is the sensor-laser detuning. The trivial time-dependence of this coupling can be removed by incorporating a Hamiltonian term $\Delta_{\xi}\hat{\xi}^{\dagger}\hat{\xi}$, which yields a simple time-independent coupling $g_{v}=\sqrt{\Gamma}$ (upon setting $\phi=\pi$), in direct correspondence to the sensor approach from Refs. [44, 45, 46].
Consequently, the cascaded master equation governing the joint dynamics of the emitter and the ancillary sensor cavity—described by their reduced density matrix, $\hat{\rho}_{\mathrm{q,1}}$— is given by [see details in Appendix A]
@@
Additionally, we have introduced an imperfect coupling factor $\varepsilon\leq 1$, accounting, e.g., for possible losses in the transmission from the source to the target [48] or to retain the scattered light from a coherently driven system [45, 46]. Therefore, $\varepsilon=1$ corresponds to perfect cascaded coupling, and $\varepsilon=0$ to the limit of no coupling.
Although the single-sensor case is relatively straightforward to describe mathematically, the extension to two or more cascaded sensors requires additional attention, since there are multiple descriptions which correspond to different physical implementations of the multi-mode measurement. In this work, we will explore the specific case in which, prior to the filtering of each mode, the output signal is physically split [66, 58] (e.g., by a balanced beam splitter in a Hanbury-Brown Twiss setup), as illustrated in Fig. 1(a). The resulting dynamics of the quantum emitter and the two frequency-filtered modes is described by the joint density matrix $\hat{\rho}_{\mathrm{q,2}}$, whose evolution obeys the cascaded master equation [see details in Appendix A]
$\frac{d\hat{\rho}_{\mathrm{q,2}}}{dt}=-i[\hat{H}+\sum_{i=1}^{2}\Delta_{i}\hat{\xi}_{i}^{\dagger}\hat{\xi}_{i},\hat{\rho}_{\mathrm{q,2}}]+\frac{\gamma}{2}\mathcal{D}[\hat{\sigma}]\hat{\rho}_{\mathrm{q,2}}\\ +\sum_{i=1}^{2}\frac{\Gamma_{i}}{2}\mathcal{D}[\hat{\xi}_{i}]\hat{\rho}_{\mathrm{q,2}}-\sqrt{\varepsilon\gamma\Gamma_{i}/2}\left\{[\hat{\xi}_{i}^{\dagger},\hat{\sigma}\hat{\rho}_{\mathrm{q,2}}]+[\hat{\rho}_{\mathrm{q,2}}\hat{\sigma}^{\dagger},\hat{\xi}_{i}]\right\}.$ (7)
In both cascaded master equations, Eq (6) and Eq. (7), $\hat{H}$ denotes the Hamiltonian of the quantum source as given in Eq. (1). The operator $\hat{\xi}_{i}$ represents the $i$th bosonic mode associated with each sensor, characterized by a sensor-laser detuning $\Delta_{i}\equiv\omega_{i}-\omega_{L}$, and linewidth $\Gamma_{i}$. To relax the notation in the two-sensor case, we have replaced the variable $\Delta_{\xi}\rightarrow\Delta_{i}$ ($i=1,2$). Throughout this work we consider identical sensor linewidths, $\Gamma_{1}=\Gamma_{2}=\Gamma$, and thus identical cascaded coupling strengths due to Eq. (5). The factor $1/\sqrt{2}$ in Eq. (7) accounts for vacuum contributions arising from signal splitting at the balanced beam splitter. In contrast, in a previous work [41], we considered the situation in a waveguide setup in which the photonic modes are digitally filtered by post-processing the signal, resulting in a different master equation from Eq. (7) that does not include the factor $1/\sqrt{2}$ in the cascaded-coupling term. As we mentioned, extensions to $N\geq 2$ sensors would require additional attention. These could be implemented using multiport beam-splitter, as in multi-copy quantum information protocols [67, 68, 69], by employing cascaded cavity arrays [39, 40] or frequency-dispersed photon detection schemes [52].
@@ II.2 Metrological strategy
where $\hat{D}(\vec{\alpha})=\prod_{i=1}^{N}\hat{D}(\alpha_{i})$ is the multi-mode displacement operator, with each $\alpha_{i}\in\mathbb{C}$ corresponding to the coherent displacement applied to the $i$th sensor, and $|\vec{n}\rangle\equiv|n_{1},\ldots,n_{N}\rangle$, with $|n_{i}\rangle$ denoting the photon number eigenstates of the $i$th sensor. We implement this strategy by performing $N$ discrete frequency-resolved displaced photon-counting measurements, with each frequency channel modeled as an ancillary single-mode bosonic cavity. We emphasize that the number of detection channels $N$ corresponds to the number of modeled frequency filters—sensors—, not to repeated experimental realizations. Formally, the measurement process is described by the tensor product of sets of operators describing the POVMs, $\hat{\Lambda}=\{\bigotimes_{i=1}^{N}\hat{\Lambda}_{n,\alpha_{i}}^{(i)}\}$, where
$\hat{\Lambda}_{n,\alpha_{i}}^{(i)}=\hat{D}^{\dagger}(\alpha_{i})|n_{i}\rangle\langle n_{i}|\hat{D}(\alpha_{i}).$ (10)
Physically, displaced photon-counting can be realized by mixing the output field of the quantum emitter with a local oscillator $\alpha(t)$ via a highly transmissive beam splitter, yielding an effective detection signal $\sqrt{\gamma}\hat{\sigma}(t)+\alpha(t)$. This mean-field engineering technique [79, 80, 81] enables us to tailor the photonic state to either isolate the quantum fluctuations or optimize the entire state for the estimation of specific atomic parameters. Standard photon-counting corresponds to the special case $\vec{\alpha}=\vec{0}$, denoted as $p(\vec{n}|\theta)\equiv p^{\vec{\alpha}=\vec{0}}(\vec{n}|\theta)$ and $F_{\theta}\equiv F^{\vec{\alpha}=\vec{0}}_{\theta}$.
Although this formulation applies generally to any number of frequency-resolved sensors, we restrict our analysis to the one‐ and two‐sensor cases. Accordingly, we evaluate $p^{\alpha}(n|\theta)$ for $N=1$ and $p^{\alpha_{1},\alpha_{2}}(n_{1},n_{2}|\theta)$ for $N=2$, as illustrated in Fig. 1(b) and (c), respectively. In this work, $\theta$ represents a Hamiltonian or a Lindblad parameter from the source master equation in Eq. (2), such as the qubit-laser detuning, the driving strength of the laser interacting with the quantum emitter, or the emitter decay rate, $\theta\in\{\Delta,\Omega,\gamma\}$.
@@
$\Delta^{2}\theta\geq\frac{1}{M\text{SNR}_{\theta}}\geq\frac{1}{MF^{\vec{\alpha}}_{\theta}}\geq\frac{1}{MI_{\theta}}.$ (12)
where $M$ is the number of repetitions, and $I_{\theta}$ denotes the quantum Fisher information (QFI), defined as the optimization of the CFI over all POVMs, or equivalently, the optimization of the SNR over all observables [17], $I_{\theta}=\text{max}_{\{\hat{\Lambda}\}}F_{\theta}=\text{max}_{\{\hat{O}\}}\text{SNR}_{\theta}$. For a general mixed state, the QFI can be expressed in closed form as [72]
$I_{\theta}=\sum_{n,m}\frac{2}{\rho_{n}+\rho_{m}}|\langle\psi_{n}|\partial_{\theta}\hat{\rho}_{\theta}|\psi_{m}\rangle|^{2},$ (13)
where $\hat{\rho}_{\theta}=\sum_{n}\rho_{n}|\psi_{n}\rangle\langle\psi_{n}|$ is the eigen-decomposition of the quantum state. Throughout this work, we use these three figures of merit—QFI, CFI, and SNR—to characterize the metrological performance of our proposed strategy. We finally give here a relevant caveat regarding the QFI. Here, the QFI is computed over the quantum state of the sensors, which captures the filtered version of the emitted light. While it bounds the precision achievable with optimal measurements on the sensor modes, it does not necessarily correspond to the absolute lower bound for $\Delta^{2}\hat{\theta}$, since it excludes information potentially available in the remaning unmeasured modes of the total radiation of the emitter. This absolute value is quantified by general expressions of the QFI as developed in Refs. [35, 24].
Figure 1(b) illustrates the application of our framework for the single-sensor case. The radiation emitted by the source is filtered by the sensor, yielding a frequency-resolved density matrix from which the conditional photon-count distribution $p(n|\theta)$ is extracted [top panel]. In this example, the distribution, obtained via standard photon-counting measurements ($\alpha=0$), forms the basis for estimating the parameter $\theta$ with a precision set by the CFI, shown by the blue solid curve. The example highlights the advantage of using the full photon-count statistics over relying only on the mean photon number. This is evident from the comparison between the CFI and the SNR of the sensor photon number (blue dashed curve). Both quantities lie below the QFI (black solid curve), indicating that neither corresponds to the optimal measurement for estimating the detuning ($\theta=\Delta$). As an illustration of the extension to the two-sensor case, Fig. 1(c) displays the joint distribution $p(n_{1},n_{2}|\theta)$ for $\alpha=0$ for the detection at two different frequencies, revealing off-diagonal structure that signals cross-correlations between detection channels, a feature that can be harnessed to enhance parameter estimation.